Topic: Discrete Mathematics and its Applications" Chapter 9:Equivalence Relations and Partial Orderings.

9. hash function H : {0,1) → {0,1) maps a bit string of length n to a bit string of length k. Hash functions are used to give a short label to a long string. The set of all collisions with a given string s defines an equivalence class for a given hash function H, that is: (a) What is the average cardinality of the equivalence classes [s]H in terms of n and k, where n > k? (b) What is the probability for any two random bit strings from s,te { 0, 1 }n that
Answer #1

Answer to (a):

The number different binary string of length n = 2n and the number different binary string of length k = 2k

So we need to map 2n strings to 2k strings.

When n > k, then each of 2k strings will be mapped on average by 2n/2k strings out of 2n strings.

Therefore, the average cardinality of the equivalence class [s]H = 2n/2k = 2n-k

Answer to (b):

The number of equivalence class = 2k

So the probability of a random string t \epsilon \{0,1\}^n to be a member of a particular equivalence class = 1/2k = 2-k

Therefore, the probability for any two random strings s,t \epsilon \{0,1\}^n such that [s]H = [t]H= 2-k

NB: Hope it helps to get your answer. Let me know any concern.